OpenStudy (anonymous):

An animal shelter has five times as many cats as dogs. Find the total number of cats if the total number of cats and dogs is 48.

OpenStudy (anonymous):

Call dogs x and cats y. so from the information you can write two equations Five times as many cats as dogs 5x=y total number of cats and dogs is 48 x+y=48 now you have two simultaneous equations 5x=y x+y=48

OpenStudy (anonymous):

Do you know how to solve now?

OpenStudy (anonymous):

i need a little more help

OpenStudy (anonymous):

Did you get stuck?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Well the first equation is already solved for y, so you can plug that expression into the other equation for y x+5x=48 now just solve for x

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

wait why is there 2 x's

OpenStudy (anonymous):

In the first equation 5x=y so since y equals 5x we can plug 5x into the second equation in place of y, so we only have one variable to solve for

OpenStudy (anonymous):

is it 8

OpenStudy (anonymous):

Yes x=8 so now that you know what x is, plug it into either equation and then solve for y

OpenStudy (anonymous):

6

OpenStudy (anonymous):

plug 8 into the second equation for x... 8+y=48 so y=?

OpenStudy (anonymous):

40

OpenStudy (anonymous):

Right so thats the number of cats

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

You're welcome